According to a 2009 Reader's Digest article, people throw away approximately 12% of what they buy at the grocery store. Assume this is the true proportion and you plan to randomly survey 229 grocery shoppers to investigate their behavior. What is the probability that the sample proportion is between 0.05 and 0.15?

Answer =______ (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Here

Population priportion =p =0.12

Sample size = n = 229

n*p = 229 * 0.12 =27.48

n*(1-p) = 229 *(1-0.12) = 201.52

Both n*p and n*(1-p) are greater than 10.

So, the sample proportion is approximately normal with

mean = \mu _{\widehat{p}} =p = 0.12

and standard deviation given by

\sigma _{\widehat{p}}=\sqrt{\frac{p(1-p)}{n}}=\sqrt{\frac{0.12(1-0.12)}{229}}\approx 0.0215

Now,

P(0.05<\widehat{p}<0.15)=P(\frac{0.05-0.12}{0.0215}<\frac{\widehat{p}-\mu _{\widehat{p}}}{\sigma _{\widehat{p}}}<\frac{0.15-0.12}{0.0215})

   =P(-3.26<Z<1.40) ( round to two decimals place)

  =P(Z<1.40)-P(Z<-3.26)

=0.9192-0.0006 ( from standard normal table)

  =0.9186

The probability that the sample proportion is between 0.05 and 0.15 is 0.9186